OpenStudy (anonymous):

NEED HELP WITH 10 QUESTIONS, I GIVE MEDALS PLEASE HELP. Solve the equation 6n + 11 = -7 + 3n for n. n equals negative two n equals negative six n equals six n equals two

OpenStudy (anonymous):

6n + 11 = -7 + 3n 6n-3n= -7-11 3n= -18 n= -18/3 n=-6

OpenStudy (anonymous):

No, that doesn't work. It's 6n+11=-7+3n 6n+4=3n 6n=3n-4 3n=4 n=1.333(on to infinity)

OpenStudy (anonymous):

n equals 2

OpenStudy (anonymous):

I know it's n=-6 because 6n+11=-7+3n -3n -3n 3n+11=-7 -11 -11 3n=-18 3n/3 -18/3 n=-6